SPECTROSCOPIC RESEARCH OF Pt + NH_{3}^{a}

\underline{JAMIE GENGLER}, TIMOTHY STEIMLE, and JINHAI CHEN, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The reaction products of laser ablated platinum with an ammonia/argon supersonic gas expansion mixture have been monitored in the 13500 cm^{−1} to 12400 cm^{−1} spectral region using laser induced fluorescence (LIF) spectroscopy. Four band features with heads near 13440 cm^{−1}, 13250 cm^{−1}, 13120 cm^{−1}, and 12435 cm^{−1} were detected. Analysis of the dispersed fluorescence suggests that the carriers of the latter three bands are Pt_{2}^{b}, PtN, and PtN. Possible electronic transition assignments for these systems based upon existing electronic structure predictions for Pt_{2}^{c} and PtN^{d} will be given. The dispersed fluorescence spectrum for the 13440 cm^{−1} band, which exhibits two progressions having ΔG_{v} of ∼ 450cm^{−1} and ∼ 650cm^{−1}, suggests that the carrier of this band is polyatomic. The excitation LIF and dispersed LIF spectra are not altered upon substitution of NH_{3} with ND_{3} suggesting that the carrier is PtN_{2} with linear asymmetric geometry^{e}’^{f}. The high resolution LIF spectrum (FWHM = 35 MHz) of the 13440 cm^{−1} band of a molecular beam sample reveals rotational structure which can be assigned to the major ^{194}Pt, ^{195}Pt, ^{196}Pt, and ^{198}Pt isotopomers. Progress on the analysis of the high resolution LIF and optical Stark measurements for the 13440 cm^{−1} band will be presented.

aFunded by DoE - Basic Energy Sciences

bJ.C. Fabbi et al, J. Chem. Phys. 115, 7543 (2001).

cK. Balsubramanian, J. Chem. Phys. 87, 6573 (1987).

dK.Y. Jung, T.C. Steimle et al, J. Chem. Phys. 102, 643 (1995).

eA. Citra et al, J. Phys. Chem. A 105, 7799 (2001).

fD.W. Green et al, J. Chem. Phys. 58, 5453 (1973).